# A reconfigurable entanglement distribution network suitable for connecting multiple ground nodes with a satellite

### v1

@@
A reconfigurable entanglement distribution network suitable for connecting multiple ground nodes with a satellite
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
Bridging the long distances between different metropolitan-scale quantum networks is currently best achieved using satellites, which can cover thousands of kilometres. The development of efficient approaches to establish connectivity between ground nodes and a satellite is therefore an essential next step. We propose a network with dual-functionality: during a brief satellite pass, the ground network is configured as a multipoint-to-point topology where all ground nodes establish entanglement with a satellite receiver. During times when this satellite is not available, the satellite up-link is rerouted via a single optical switch to the ground nodes, and the network is configured as a pair-wise ground network. We numerically simulate a pulsed hyper-entangled photon source and study the performance of the proposed network configurations for quantum key distribution. We find favourable resource overheads in the case that the satellite receiver exploits time-multiplexing whereas the ground nodes utilize frequency-multiplexing. Our results show high and scalable key rates can be achieved in both configurations. The scalability, simple reconfigurability, and easy integration with fibre networks make this architecture a promising candidate for quantum communication of many ground nodes and a satellite thus paving the way towards interconnection of ground nodes at a global scale.
@@ 1 Introduction
A primary objective of quantum communication is the distribution of entanglement between distant parties [1]. This serves as the foundation for creating the quantum internet: a quantum network that enables revolutionary applications, including quantum key distribution (QKD) [2, 3], clock synchronization [4, 5], distributed quantum computing [6, 7], and long-baseline interferometry [8]. In recent years, significant research efforts have been done in expanding the scalability of quantum networks beyond two communicating parties. To this extent, multiple network architectures have been proposed, notably trusted-node networks which consists of a mesh of point-to-point links, where each node acts as a trusted relay to establish a complete two-party communication set-up [9, 10, 11, 12, 13]. While such a scheme is compatible with heterogeneous quantum links and can be used to increase the communication distance arbitrarily, it comes with a significant security cost as one must ensure the privacy of every relay site. Alternatively, active switch quantum networks [14, 15, 16, 17, 18] allow for dynamically reconfigurable networks yet are vulnerable to device failure. Furthermore, the distribution rate is limited by the switch’s operating speed. Hence, an all-passive solution is preferred where all users can be connected simultaneously.
@@
Thus far, two main approaches have been employed to create fully-connected quantum networks relying either on high-dimensional entanglement or bi-partite entanglement and multiplexing [19, 9, 20, 21, 22, 23]. In the multipartite method, each user pair shares a different subspace of the overall network Hilbert space, which is used to generate their keys. However, adding or subtracting users from the network requires changing the dimensionality of the state produced by the source making this approach unscalable. On the other hand, multiplexing exploits the inherent hypercorrelations produced by the entangled photon source to deterministically separated correlated photon pairs to different users. This architecture is easily scalable as the number of users in the network can be modified without any change to the photon source. Nevertheless for an $n$-node network to be fully connected, $\mathcal{O}(n^{2})$ multiplexed channels are required. Consequently, there remains an important hardware requirement as each channel requires a decoder and detector module to be fully distinguishable. Otherwise an increase in accidental counts will deteriorate the performance of the network as was shown in [9, 22]. Recently, [24, 25, 26, 27] utilized beamsplitters in addition to multiplexing to increase the scalability of their network, but at the expense of the key generation rate. An interesting solution proposed by [9], is to use a pulsed source to recover the channel distinguishability despite limited detector resources. This implementation requires an additional gating signal to be sent to each user, and results in noise reduction by a factor equal to the gating duty cycle. However, this scheme is unsuitable for mobile nodes as the delays to every other node would need to be continuously compensated by each user thus restricting the reach of these quantum networks to metropolitan areas. Continued extensive research is being conducted on fibre-based quantum repeaters [28, 29, 30] and satellite-based quantum links [31, 32, 33, 34] to enable the establishment of a global-scale network. While fibre-based links are limited in distance by exponential losses, the use of a free-space link to an orbiting satellite reduces the loss scaling from exponential to quadratic, consequently extending the reach of individual quantum links [35].
@@ 2 Method
The proposed network architecture is depicted in Fig. 1. It allows two reconfigurable quantum metropolitan-area networks (QMAN), for example in Waterloo and Montreal, Canada, to be connected by ground-to-satellite quantum links. As the satellite passes over Waterloo in Fig. 1b, the network adopts a multipoint-to-point topology from the $N$ ground nodes to the satellite. Consequently allowing for a substantive increase of the key generation during the satellite pass via multiplexing. Meanwhile, the Montreal network in Fig. 1c is arranged into a pair-wise configuration where every user is connected to each other. As the satellite travels out of reach of Waterloo and within reach of Montreal, each network reconfigures itself into the opposite topology to optimize its functionality throughout the day.
This rearrangement is enabled by the hybrid quantum network design in Fig 2. For a satellite up-link, the idler is kept on the ground and distributed to the network users. During the satellite pass, the signal is sent through a frequency-to-time encoder $\uptau(\lambda)$ to assign a time delay to each wavelength channel. This mapping can be accomplished by wavelength-dependent group delay dispersion e.g. in a chirped fibre Bragg grating [36, 39, 40] (CFBG) or by using a fibre Bragg grating (FBG) array with suitable spacing between each FBG to obtain the desired time delay [37]. Outside of the satellite pass, the flip mirror redirects the signal to the ground nodes.
The minimal duration of the channel time separation for the frequency-to-time mapping is determined by the timing jitter of the satellite single-photon detector and readout electronics which is typically on the order of $\sim 130$ ps for Si-APDs [41, 42, 43, 44]. Thus for a 80 MHz pump repetition rate, the satellite network configuration is limited to a maximum of $\sim 95$ wavelength-time channels within the 12.5 ns pulse train. To ensure the channel distinguishability, the GDD must then be greater than $\sim$ 325 ps/nm. For a given device length, the GDD and the CFBG’s reflection bandwidth are inversely proportional: a larger dispersion requires higher reflectivities resulting in a narrower resonance and bandwidth [45, 46]. However, this limitation can be circumvented by using longer CFBGs or by using a CFBG array i.e. inscribing or splicing multiple CFBGs with adjacent reflection windows in succession [36].
@@
3 Performance estimation
$\hat{E}^{(\text{click})}=\sum_{m=0}^{\infty}[1-D(0)(1-\eta)^{m}]\outerproduct{m}{m}.$ | . (2)
and
$\hat{E}^{(n)}=\sum_{m=0}^{\infty}\sum_{k=0}^{n}D(k)\binom{n-k+m}{m}\eta^{n-k}(1-\eta)^{m}\outerproduct{n+m-k}{n+m-k},$ (3)
where $\eta$ is the detector efficiency and $D(k)$ the dark count distribution. For the satellite pass configuration, we assume the satellite detector to be a Si-APD [53] with a detector dead time of 1 $\mu$s whereas we assume the ground detector to be an SNSPD with a dead time of 10 ns [54, 55]. Due to the dead time, the photon count rate is no longer linearly proportional to the impinging photon flux but is given by:
@@ 3 Performance estimation
At low ground link loss ($<10$ dB) in Fig. 3(a), detector saturation at the satellite node partially limits the scaling performance. However, in the higher loss regime one obtains a linear key rate increase whilst maintaining the same QBER. This scaling stems from the deterministic separation of the multiplexed channels resulting in independent communication channels with the same signal-to-noise ratio [9]. To achieve this deterministic scenario, the frequency-to-time $\uptau(\lambda)$ plays a crucial role. This is particularly apparent, in Fig. 3(b) where we only consider wavelength multiplexing. In other words, the idler photons are demultiplexed on the ground but the signal photons are recorded at the satellite detector without any channel discrimination. Surprisingly, demultiplexing solely the idler at the ground stations provides little advantage and in fact results in a maximal tolerable loss decrease compared to a single channel. This is due to an increase in accidental coincidences resulting in a loss of fidelity caused by detecting several channels on the single satellite detector [22]. Therefore, the signal-to-noise ratio decreases proportionally to the number of users added. Note that for every channel link setting, the squeezing parameter ($\chi$) was numerically optimized to maximize the SKR. To compensate for the increase in accidental counts at higher loss, the photon pair production was decreased proportionally to the number of channels considered. At low loss, this one-sided multiplexing approach yields a slight key rate increase due to detector load management bypassing the dead time of the detector. Nonetheless, for realistic loss scenarios detector saturation has negligible significance. We investigate in Fig. 4 the use of PNR detectors at the ground user nodes. Interestingly, we find that PNR detectors offer no significant key rate advantage compared to bucket detectors. This can be readily seen in the high loss regime from considering the series expansion of $\hat{E}^{(1)}$ for $D(1)\ll 1,\eta\to 0$ at which point one finds that $\hat{E}^{(1)}\to\hat{E}^{(\text{click})}$. A detailed analysis of this phenomenon is in preparation [57].
@@
Figure 5: The spectrum of the signal (a) and the idler (b) were calculated from the Sellmeier equations for PPMgOLN assuming a 2 nm pump bandwidth centered at 521.4 nm and a 15 mm crystal [[60]]. Wavelength correlated pairs $\{\lambda_{i},\Lambda_{i}\}$ are denoted by the same color in both spectrums and were chosen in accordance with Fig. 1. The idler channels match the ITU DWDM 100 GHz grid.
@@ 4 Conclusion
We proposed and analyzed an architecture suitable for serving both metropolitan networks or ground-satellite links, reconfigurable with minimal modifications. Our approach is uniquely designed to optimize the efficiency of satellite up-link scenarios such as QEYSSat, where the satellite receiver is only available for a few minutes every 24 hours. During the satellite pass, the secure key rate is improved through time and frequency multiplexing with no additional hardware requirements on the satellite. Additionally, the network’s utility outside of the satellite pass is optimized by rerouting the signal to form a ground-based metropolitan network. The satellite can be exploited as a trusted node or entanglement swapping station to interconnect two metropolitan areas and extend the reach of quantum networks to a global scale. The scalability and simplicity render our network architecture a good candidate for commercial quantum communication networks.
@@
[…]
@@ A.1 Squashing Model
def measure_2folds_4modes_squashing_optimized(dim,$\chi$,proj,proj2,$\tau$):
@@
       H_sq = 1j*$\chi$*(tensor(create(dim),create(dim))+tensor(destroy(dim),destroy(dim)))
@@
       out_i = reshape(transpose(reshape(psi_i.full(), (dim,dim,dim,-1)),(0,3,2,1)),(dim*dim*dim*dim,-1))
@@
                               det_exp[i][j][k][l] = abs(expect(tensor(proj[i],proj[j],proj2[k],proj2[l]),psi_i))
       HH $=\frac{(\rm det\_exp[0][1][0][1]+0.5*(det\_exp[0][0][0][1]+det\_exp[0][1][0][0])+0.25*det\_exp[0][0][0][0])}{\rm(1+\tau*(det\_exp[0][1][0][1]+0.5*(det\_exp[0][0][0][1]+det\_exp[0][1][0][0])+0.25*det\_exp[0][0][0][0]))}$
       VV $=\frac{\rm(det\_exp[1][0][1][0]+0.5*(det\_exp[0][0][1][0]+det\_exp[1][0][0][0])+0.25*det\_exp[0][0][0][0])}{\rm(1+\tau*(det\_exp[1][0][1][0]+0.5*(det\_exp[0][0][1][0]+det\_exp[1][0][0][0])+0.25*det\_exp[0][0][0][0]))}$
       HV$=\frac{\rm(det\_exp[0][1][1][0]+0.5*(det\_exp[0][0][1][0]+det\_exp[0][1][0][0])+0.25*det\_exp[0][0][0][0])}{\rm(1+\tau*(det\_exp[0][1][1][0]+0.5*(det\_exp[0][0][1][0]+det\_exp[0][1][0][0])+0.25*det\_exp[0][0][0][0]))}$
       VH$=\frac{\rm(det\_exp[1][0][0][1]+0.5*(det\_exp[0][0][0][1]+det\_exp[1][0][0][0])+0.25*det\_exp[0][0][0][0])}{\rm(1+\tau*(det\_exp[1][0][0][1]+0.5*(det\_exp[0][0][0][1]+det\_exp[1][0][0][0])+0.25*det\_exp[0][0][0][0]))}$
@@ A.2 Secure key rate calculation
def simqkdentanglement($\chi$,loss_a,loss_b,n_factor_a,n_factor_b,dim,N):
@@
            a_det = BucketDetector_realistic_detector(N,loss_a,n_factor_a)
@@
             probs2f_i= measure_2folds_4modes_squashing_optimized(N,$\chi$,a_det,b_det,$\tau$)

### v2

@@
Abstract
Satellite-based quantum communication channels are important for ultra-long distances. Given the short duration of a satellite pass, it can be challenging to efficiently connect multiple users of a city-wide network while the satellite is passing over that area. We propose a network with dual-functionality: during a brief satellite pass, the ground network is configured as a multipoint-to-point topology where all ground nodes establish entanglement with a satellite receiver. During times when this satellite is not available, the satellite up-link is rerouted via a single optical switch to the ground nodes, and the network is configured as a pair-wise ground network. We numerically simulate a pulsed hyper-entangled photon source and study the performance of the proposed network configurations for quantum key distribution. We find favourable scaling in the case that the satellite receiver exploits time-multiplexing whereas the ground nodes utilize frequency-multiplexing. The scalability, simple reconfigurability, and easy integration with fibre networks make this architecture a promising candidate for quantum communication of many ground nodes and a satellite, thus paving the way towards interconnection of ground nodes at a global scale.
keywords:
Quantum networks, quantum communication, quantum key distribution, satellite-based quantum information
@@ 1 Introduction
A primary objective of quantum communication is the distribution of entanglement between distant parties [1]. This serves as the foundation for creating the quantum internet: a quantum network that enables revolutionary applications, including quantum key distribution (QKD) [2, 3], clock synchronization [4, 5], distributed quantum computing [6, 7], and long-baseline interferometry [8]. In recent years, significant research efforts have been done in expanding the scalability of quantum networks beyond two communicating parties. To this extent, multiple network architectures have been proposed, notably trusted-node networks which consist of a mesh of point-to-point links, where each node acts as a trusted relay to establish a complete two-party communication set-up [9, 10, 11, 12, 13]. While such a scheme is compatible with heterogeneous quantum links and can be used to increase the communication distance arbitrarily, it comes with a significant security cost as one must ensure the privacy of every relay site. Alternatively, active switch quantum networks [14, 15, 16, 17, 18] allow for dynamically reconfigurable networks, yet are vulnerable to device failure. Furthermore, the distribution rate is limited by the switch’s operating speed. Hence, an all-passive solution is preferred where all users can be connected simultaneously.
@@
Two main approaches have previously been considered to create fully-connected quantum networks relying either on high-dimensional entanglement or bi-partite entanglement and multiplexing [19, 9, 20, 21, 22, 23, 24]. In the multipartite method, each user pair shares a different subspace of the overall network Hilbert space, which is used to generate their keys. However, adding or subtracting users from the network requires changing the dimensionality of the state produced by the source making this approach unscalable. On the other hand, multiplexing exploits the inherent hypercorrelations produced by the entangled photon source to deterministically separate correlated photon pairs to different users. This architecture is easily scalable as the number of users in the network can be modified without any change to the photon source. Nevertheless for an $n$-node network to be fully connected, $\mathcal{O}(n^{2})$ multiplexed channels are required. Consequently, there remains an important hardware requirement as each channel requires a decoder and detector module to be fully distinguishable. Otherwise an increase in accidental counts will deteriorate the performance of the network as was shown in [9, 22]. Recently, [25, 26, 27, 28] utilized beamsplitters in addition to multiplexing to increase the scalability of their network, but at the expense of the key generation rate. An interesting solution proposed by [9], is to use a pulsed source to recover the channel distinguishability despite limited detector resources. This implementation requires an additional gating signal to be sent to each user, and results in noise reduction by a factor equal to the gating duty cycle. However, this scheme is unsuitable for mobile nodes as the delays to every other node would need to be continuously compensated by each user thus restricting the reach of these quantum networks to metropolitan areas. Continued extensive research is being conducted on fibre-based quantum repeaters [29, 30, 31] and satellite-based quantum links [32, 33, 34, 35] to enable the establishment of a global-scale network. While fibre-based links are limited in distance by exponential losses, the use of a free-space link to an orbiting satellite reduces the loss scaling from exponential to quadratic, consequently extending the reach of individual quantum links [36].
@@ 2 Method
The proposed network architecture is depicted in Fig. 1. It allows two reconfigurable quantum metropolitan-area networks (QMAN), for example in Waterloo and Montreal, Canada, to be connected by ground-to-satellite quantum links. As the satellite passes over Waterloo in Fig. 1b, the network adopts a multipoint-to-point topology from the $N$ ground nodes to the satellite. Consequently allowing for a substantive increase of the key generation during the satellite pass via multiplexing. Meanwhile, the Montreal network in Fig. 1c is arranged into a pair-wise configuration where every user is connected to each other. As the satellite travels out of reach of Waterloo and within reach of Montreal, each network reconfigures itself into the opposite topology to optimize its functionality throughout the day. This rearrangement is enabled by the hybrid quantum network design in Fig 2. For a satellite up-link, the idler is kept on the ground and distributed to the network users. During the satellite pass, the signal is sent through a frequency-to-time encoder $\uptau(\lambda)$ to assign a time delay to each wavelength channel. This mapping can be accomplished by wavelength-dependent group delay dispersion e.g. in a chirped fibre Bragg grating [37, 40, 41] (CFBG) or by using a fibre Bragg grating (FBG) array with suitable spacing between each FBG to obtain the desired time delay [38]. Outside of the satellite pass, the flip mirror redirects the signal to the ground nodes.
The minimal duration of the channel time separation for the frequency-to-time mapping is determined by the timing jitter of the satellite single-photon detector and readout electronics which is typically on the order of $\sim 130$ ps for Si-APDs [42, 43, 44, 45]. Thus for a 80 MHz pump repetition rate, the satellite network configuration is limited to a maximum of $\sim 95$ frequency-time channels within the 12.5 ns pulse train. To ensure the channel distinguishability, the GDD must then be greater than $\sim$ 325 ps/nm. For a given device length, the GDD and the CFBG’s reflection bandwidth are inversely proportional: a larger dispersion requires higher reflectivities resulting in a narrower resonance and bandwidth [46, 47]. However, this limitation can be circumvented by using longer CFBGs or by using a CFBG array i.e. inscribing or splicing multiple CFBGs with adjacent reflection windows in succession [37].
@@
3 Performance estimation
$\hat{E}^{(\text{click})}=\sum_{m=0}^{\infty}[1-D(0)(1-\eta)^{m}]\ket{m}\bra{m}.$ | . (2)
and
$\hat{E}^{(n)}=\sum_{m=0}^{\infty}\sum_{k=0}^{n}D(k)\binom{n-k+m}{m}\eta^{n-k}(1-\eta)^{m}\ket{n+m-k}\bra{n+m-k},$ (3)
where $\eta$ is the detector efficiency and $D(k)$ the dark count distribution. For the satellite pass configuration, we assume the satellite detector to be a Si-APD [54] with a detector dead time of 1 $\mu$s whereas we assume the ground detector to be an SNSPD with a dead time of 10 ns [55, 56].
@@ 3 Performance estimation
Due to the dead time, the photon count rate is no longer linearly proportional to the impinging photon flux but is given by:
@@
At low ground link loss ($<10$ dB) in Fig. 3(a), detector saturation at the satellite node partially limits the scaling performance. However, in the higher loss regime one obtains a linear key rate increase whilst maintaining the same QBER. This scaling stems from the deterministic separation of the multiplexed channels resulting in independent communication channels with the same signal-to-noise ratio [9]. To achieve this deterministic scenario, the frequency-to-time mapping $\uptau(\lambda)$ plays a crucial role. This is particularly apparent, in Fig. 3(b) where we only consider wavelength multiplexing. In other words, the idler photons are demultiplexed on the ground but the signal photons are recorded at the satellite detector without any channel discrimination. Surprisingly, demultiplexing solely the idler at the ground stations provides little advantage and in fact results in a maximal tolerable loss decrease compared to a single channel. This is due to an increase in accidental coincidences resulting in a loss of fidelity caused by detecting several channels on the single satellite detector [22]. Therefore, the signal-to-noise ratio decreases proportionally to the number of users added. Note that for every channel link setting, the squeezing parameter ($\chi$) was numerically optimized to maximize the SKR. To compensate for the increase in accidental counts at higher loss, the photon pair production was decreased proportionally to the number of channels considered. At low loss, this one-sided multiplexing approach yields a slight key rate increase due to detector load management bypassing the dead time of the detector. Nonetheless, for realistic loss scenarios detector saturation has negligible significance. We investigate in Fig. 4 the use of PNR detectors at the ground user nodes. Interestingly, we find that PNR detectors offer no significant key rate advantage compared to bucket detectors. This can be readily seen in the high loss regime from considering the series expansion of $\hat{E}^{(1)}$ for $D(1)\ll 1,\eta\to 0$ at which point one finds that $\hat{E}^{(1)}\to\hat{E}^{(\text{click})}$. A detailed analysis of this phenomenon is in preparation [58].
@@
Figure 5: The spectrum of the signal (a) and the idler (b) were calculated from the Sellmeier equations for PPMgOLN assuming a 2 nm pump bandwidth centered at 521.4 nm and a 15 mm crystal [61]. Wavelength correlated pairs $\{\lambda_{i},\Lambda_{i}\}$ are denoted by the same color in both spectrums and were chosen in accordance with Fig. 1. The idler channels match the ITU DWDM 100 GHz grid.
@@ 4 Conclusion
Declarations
Funding
This work is supported by the High Throughput and Secure Networks (HTSN) challenge by the National Research Council Canada, project HTSN-630. The authors further acknowlege support from the Natural Sciences and Engineering Research Council of Canada (NSERC), the Canadian Foundation for Innovation (CFI), the Ontario Research Fund (ORF). SV would like to thank the NSERC PGS-D for personal funding.
Conflict of interest/Competing interests
The authors declare no conflicts of interest related to this article.
Code availability
The simulation code used in this paper is available in the supplementary materials, Ref. [63].
@@ A.1 Squashing Model
def measure_2folds_4modes_squashing_optimized(dim,chi,proj,proj2,tau):
@@
[…]
@@
    out_i = reshape(transpose(reshape(psi_i.full(), (dim,dim,dim,-1)),(0,3,2,1)),
    (dim*dim*dim*dim,-1))
@@
                 det_exp[i][j][k][l]} =
                 abs(expect(tensor(proj[i],proj[j],proj2[k],proj2[l]),psi_i))
    HH=(det_exp[0][1][0][1]+0.5*(det_exp[0][0][0][1]+det_exp[0][1][0][0])
    +0.25*det_exp[0][0][0][0])/(1+ tau*(det_exp[0][1][0][1]
    +0.5*(det_exp[0][0][0][1]+det_exp[0][1][0][0])+0.25*det_exp[0][0][0][0]))
    VV=(det_exp[1][0][1][0]+0.5*(det\_exp[0][0][1][0]+det_exp[1][0][0][0])
[…]
    +0.5*(det_exp[0][0][1][0]+det_exp[1][0][0][0])+0.25*det_exp[0][0][0][0]))
    HV=(det_exp[0][1][1][0]+0.5*(det_exp[0][0][1][0]+det_exp[0][1][0][0])
[…]
    +0.5*(det_exp[0][0][1][0]+det_exp[0][1][0][0])+0.25*det_exp[0][0][0][0]))
    VH=(det_exp[1][0][0][1]+0.5*(det_exp[0][0][0][1]+det_exp[1][0][0][0])
[…]
    +0.5*(det_exp[0][0][0][1]+det_exp[1][0][0][0])+0.25*det_exp[0][0][0][0]))
@@ A.2 Secure key rate calculation
simqkdentanglement(chi,loss_a,loss_b,n_factor_a,n_factor_b,dim,N}):
@@
        BucketDetector_realistic_detector(N,loss_a,n_factor_a)}
@@
        probs2f_i= measure_2folds_4modes_squashing_optimized(N,chi,a_det,b_det,tau)}